OpenStudy (anonymous):

what is the slope of a line parallel to the line with equation: 5x-y=11

OpenStudy (anonymous):

A line parallel will have the same slope, so just get it into y=mx+b form, identify the slope (m), and you have your answer. Let me know if you'd like to see it fully worked out.

OpenStudy (anonymous):

can you please

OpenStudy (anonymous):

Sure! First, you want the y on the left side so subtract 5x from both sides to get: -y=-5x+11 Then multiply everything by -1 so you get positive y on the left: y=5x-11 Now it is in y=mx+b form, where m is the slope. So: m=5 The slope of a line parallel to this one will have the same m value, so 5.

OpenStudy (anonymous):

thanks